OpenStudy (anonymous):

William needs 180 mL of 2 M H2SO4, how much 18 M H2SO4 does he need to use? Patrick has 80 mL of 12 M HCl, how much 1.25 M solution can he make? I feel like the only thing missing for both problems is the volume... but I am not sure. Would I be using the dilution formula? M1VI=M2V2?

OpenStudy (maddyx):

Dont use the dilution formula.....use the law of equivalents (i think) ... because you are not "making up" the volume....use N1V1=N2V2....convert te molarity to normality by diving M/(basicity of acid) (a) N1 = M/2 = 2/2 = 1 N So, 1 * 180 = (18/2) * V2 Therefore, V2 = 20 ml (b) Similarly this should be coming 768 ml. Hope it is correct. do let me know!

OpenStudy (anonymous):

Use the dilution formula M1V1 = M2V2 180(2) = 18(x), solve for x to get the correct answer in mL 80(12) = 1.25(x), solve for x to get the correct answer in mL